Hey guys i was wondering if someone could help me out with this word problem, thank you.
An amphitheater has 22 seats in the first row and 40 rows in all. Each successive row has one additional seat. How many seats are in the amphitheater?

Answer by ewatrrr(24785)

considering this as an Arithmetic Sequence: d = 1, a1 = 22
.......
a40 = 22 + 39(1) = 61 seats
Sum40 = 40/2(22 + 61) = 20(83) = 1660 seats in all.